OpenStudy (anonymous):

Two forces with magnitudes of 150 and 100 pounds act on an object at angles of 40° and 170° respectively. Find the direction and magnitude of the resultant force. Round to two decimal places in all intermediate steps and in your final answer.

OpenStudy (anonymous):

This can not be solve without making a figure. It is a self explaining Ref: Fig: Note : 10 degrees = 180° - 170 (@100lb) = 10° (180 is straight line) Sin 10° = D/100 D = 100 Sin 10 D = 17.36 lbs Sin 40 = E/50 E = 40 Sin 40 E = 32.14 lbs Cos 40 = G/50 G = 50 Cos 40 G = 38.3 lbs Cos 10 ={G + I)/100 (G + I ) = 100 Cos 10 (G + I ) = 98.48 lb I = (G + I) – G I = 98.48 – 38.3 I = 60.18 lbs Tan w = (E + D)/I Tan w = ( 32.14 + 17.36) / 60.18 Tan w = 0.823 Direction: =======================================... W= 39.44 ° ◄ this is made from resultant Force & the Horizontal =======================================... Sin w = (E+D)/ Ft Ft = (E+D)/Sin w Ft = (32.14 + 17.36) Sin 39.44 Ft = (32.14 + 17.36) 0.64 =======================================... Ft ≈ 31.68 lbs ◄ Ans (Resultant Force:) =======================================... Hope this helps Remember that Jesus loves you. Know Him in His word in the Bible. God Bless

OpenStudy (anonymous):

here is the figure(-: http://prntscr.com/3rvlil

OpenStudy (anonymous):

Thank you so much for the thorough explanation! I wish you clear skies!